# Large Language Models for Software Engineering: A Systematic Literature Review

### v5

@@ 4. RQ2: How are SE-related datasets collected, preprocessed, and used in LLMs?
Data plays a crucial role in the model training phase (Sun et al., 2022). First, data is collected to obtain diversity and richness to ensure that the model can cope with different scenarios and situations. Second, data is classified to clarify the training objectives of the model and avoid confusion and misinformation. The preprocessing of data is indispensable to clean and transform the data to improve its quality. Finally, data is formatted into a structure suitable for model processing, allowing the LLM to effectively learn the data’s features and patterns. We analyze the reported processes of data collection, data classification, data preprocessing, and data representation in our selected primary studies on LLM4SE.
@@ 4.1. How are the datasets for training LLMs sourced?
Data is an indispensable and critical factor in the training of LLMs, which determines the generalization ability, effectiveness, and performance of the models (Sun et al., 2022). Adequate, high-quality, and diverse data is critical to allow models to fully learn features and patterns, optimize parameters, and ensure reliability in validation and testing. We first investigate the methods used to obtain the dataset. By analyzing the methods of data collection, we divided the data sources into four categories: open-source datasets, collected datasets, constructed datasets, and industrial datasets. Open-source datasets (Chen et al., 2023c; Khakhar et al., 2023; Wang et al., 2023d; Zeng et al., 2022) refer to publicly accessible collections of data that are often disseminated through open-source platforms or repositories. For example, datasets like HumanEval (Chen et al., 2021b), which consists of 164 manually crafted Python problems, each accompanied by its respective unit tests. The open-source nature of these datasets ensures their credibility and allows for community-driven updates, making them a reliable resource for academic research. Collected datasets (Huang et al., 2018; Tian et al., 2023b; Sghaier and Sahraoui, 2023; Mastropaolo et al., 2022b) are those that researchers compile directly from a multitude of sources, including but not limited to, major websites, forums, blogs, and social media platforms. For instance, researchers (Chan et al., 2023; Salza et al., 2022; Weyssow et al., 2023; Yang et al., 2022c) often scrape data from Stack Overflow (Overflow, 2023) threads or GitHub (Github, 2023) issue comments to create a dataset tailored to their specific research questions. Constructed datasets (Ezzini et al., 2022; Koide et al., 2023; Kang et al., 2022; Zhang et al., 2022a) are specialized datasets that researchers create by modifying or augmenting collected datasets to better align with their specific research objectives. These modifications can be carried out through manual or semi-automatic methods and may include the generation of domain-specific test sets, annotated datasets, or synthetic data. For example, researchers often take a collected dataset of code snippets and manually annotate them with bug types to create a constructed dataset for studying automated program repair techniques (Fan et al., 2023a; Jin et al., 2023; Wu et al., 2023a). Industrial datasets (Alhamed and Storer, 2022; Moharil and Sharma, 2022; Wang et al., 2020c) are those obtained from commercial or industrial entities and often contain proprietary business data, user behavior logs, and other sensitive information. These datasets are particularly valuable for research that aims to address real-world business scenarios. However, the acquisition of such datasets is often complicated by issues related to business confidentiality and data privacy. For example, in a collaborative effort with China Merchants Bank (CMB), Wang et al. (Wang et al., 2020c) were able to access 21 projects from CMB’s repositories. Access to such data would likely require non-disclosure agreements and other legal safeguards to protect business interests. Each of these dataset types offers unique advantages and challenges, and the choice between them should be guided by the specific requirements and constraints of the research project at hand.
@@
Fig. 6 shows the collection strategies of LLM-related datasets. As can be seen from the data in the figure, 127 studies used open-source datasets for training large models. One of the main reasons for using open-source datasets in LLM training is their authenticity and credibility. Open-source datasets usually contain real-world data collected from various sources (such as relevant studies that have been conducted), which makes them highly reliable and representative of real-world scenarios. This helps LLMs learn from real examples to better understand real-world applications and improve their performance. Second, since LLMs are a topic that has just recently emerged, a lack of suitable training sets does exist. Therefore, researchers often collect data from sites such as Stack Overflow and GitHub and build datasets to make the data more composite for SE tasks. Out of the 229 papers we studied, we found that only four of these studies were using industrial datasets. This suggests a potential misalignment between the properties of datasets used in academic research and those encountered in real-world industrial contexts. This divergence underscores the need for future research to investigate industrial datasets, thereby ensuring that LLMs are applicable and robust across both academic and industrial scenarios.
@@ 4.2. What types of SE datasets have been used in existing LLM4SE studies?
Category Data type # Studies Total References
Code-based datasets Source code 44 70 (Ahmed et al., 2023) (Arakelyan et al., 2023) (Bui et al., 2023) (Cassano et al., 2023) (Charalambous et al., 2023) (Chen et al., 2022) (Ciniselli et al., 2021) (Gao et al., 2023) (Geng et al., 2024) (Gupta et al., 2023) (Hu et al., 2023) (Huang et al., 2023d) (Izadi et al., 2022) (Jesse et al., 2022) (Kanade et al., 2020) (Lahiri et al., 2022) (Li et al., 2021) (Li et al., 2023d) (Liu et al., 2020) (Ma et al., 2023a) (Mastropaolo et al., 2022b) (Murali et al., 2023) (Nguyen et al., 2023) (Ochs et al., 2023) (Pan et al., 2023a) (Prenner and Robbes, 2021) (Saberi et al., 2023a) (Sharma et al., 2022) (Shen et al., 2022) (Shi et al., 2022) (Siddiq et al., 2023b) (Tang et al., 2023b) (Wan et al., 2022b) (Wang et al., 2023c) (Weyssow et al., 2023) (Xie et al., 2023) (Wong et al., 2023) (Yang et al., 2023c) (Yu et al., 2023) (Zeng et al., 2022) (Zhang et al., 2023d) (Zhang et al., 2023c) (Zheng et al., 2023a) (Zhu et al., 2023)
Bugs 4 (Kang et al., 2023) (Khanfir et al., 2023) (Xia et al., 2022) (Xia and Zhang, 2023a)
Patches 4 (Le-Cong et al., 2023) (Tian et al., 2020) (Tian et al., 2023a) (Zhang et al., 2023a)
Code changes 4 (Ghadhab et al., 2021) (Li et al., 2022f) (Paul et al., 2023a) (Zhang et al., 2023f)
Test suites/cases 3 (Fan et al., 2023a) (Jana et al., 2023) (Zhang et al., 2023g)
Error code 3 (Huang et al., 2023c) (Paul et al., 2023b) (Widjojo and Treude, 2023)
Vulnerable source code 2 (Chan et al., 2023) (Chen et al., 2023a)
Bug-fix pairs 2 (Fakhoury et al., 2023) (Yuan et al., 2022)
Labeled clone pairs 1 (Dou et al., 2023)
Buggy programs 1 (Cao et al., 2023)
Packages 1 (Schäfer et al., 2023)
Flaky test cases 1 (Fatima et al., 2022)
Text-based datasets Prompts 28 104 (Cassano et al., 2023) (Dinh et al., 2023) (Dong et al., 2023) (Du et al., 2023) (Gilbert et al., 2023) (Jiang et al., 2023b) (Jiang et al., 2023a) (Ke et al., 2023) (Khakhar et al., 2023) (Li et al., 2022d) (Li et al., 2023c) (Luo et al., 2023) (Scao et al., 2022) (Siddiq et al., 2023a) (Singla, 2023) (Tan et al., 2023) (Tang et al., 2023a) (Tarassow, 2023) (Tihanyi et al., 2023) (Tu et al., 2023) (Vikram et al., 2023) (Wang et al., 2023d) (Wang et al., 2023b) (Wang et al., 2022b) (White et al., 2023a) (Xu et al., 2023) (Ye et al., 2023) (Zhao et al., 2023a)
Programming problems 14 (Chen et al., 2021b) (Chen et al., 2023c) (Hendrycks et al., 2021) (Jones and Steinhardt, 2022) (Kuznia et al., 2022) (Li et al., 2023b) (Li et al., 2023e) (Sadik et al., 2023) (Sakib et al., 2023) (Sobania et al., 2023) (Tian et al., 2023b) (Wang et al., 2023e) (Zan et al., 2022b) (Zan et al., 2023a)
SO (i.e., Stack Overflow) posts 9 (Biswas et al., 2020) (He et al., 2023) (He et al., 2022) (Helmeczi et al., 2023) (Huang et al., 2023b) (Kou et al., 2023a) (Rahmani et al., 2023) (Wei et al., 2022a) (Zhang et al., 2020b)
Bug reports 9 (Ciborowska and Damevski, 2023) (Feng and Chen, 2023) (Gomes et al., 2023) (Helmeczi et al., 2023) (Isotani et al., 2021) (Kang et al., 2022) (Lee et al., 2022) (Luitel et al., 2023) (Mukherjee and Hellendoorn, 2023)
Programming tasks (and solutions) 7 (Dibia et al., 2022) (Döderlein et al., 2022) (Fan et al., 2023a) (Khan et al., 2023) (Kou et al., 2023b) (Nijkamp et al., 2023) (Saieva et al., 2023)
Requirements documentation 6 (Ezzini et al., 2022) (Hey et al., 2020) (Kolthoff et al., 2023) (Luo et al., 2022) (Moharil and Sharma, 2023) (Wang et al., 2020c)
APIs/API documentation 5 (Deng et al., 2023c) (Khan et al., 2021) (Patil et al., 2023) (Yang et al., 2022c) (Zan et al., 2022a)
Q&A pairs 5 (Qian et al., 2023) (Schlag et al., 2023) (Von der Mosel et al., 2022) (Wang et al., 2023d) (Zhou et al., 2023a)
Vulnerability descriptions 4 (Alqarni and Azim, 2022) (Pearce et al., 2021) (Thapa et al., 2022) (Wu et al., 2023a)
Bug reports with changesets 4 (Ciborowska and Damevski, 2022) (Deng et al., 2023c) (Helmeczi et al., 2023) (Jin et al., 2023)
Methods 3 (Mastropaolo et al., 2021a) (Mastropaolo et al., 2022b) (Yuan et al., 2023b)
Code comments 2 (Prenner and Robbes, 2021) (Xu et al., 2022)
Project issues 2 (Fu and Tantithamthavorn, 2022) (Sun et al., 2023)
Software specifications 1 (Mandal et al., 2023)
Dockerfiles 1 (Henkel et al., 2021)
Semantic merge conflicts 1 (Zhang et al., 2022b)
Site text 1 (Koide et al., 2023)
User intents 1 (Jain et al., 2022)
User reviews 1 (Wang et al., 2022c)
Graph-based datasets GUI Images 1 1 (Kolthoff et al., 2023)
Software repository- Issues and commits 3 5 (Alhamed and Storer, 2022) (Lin et al., 2021) (Zhang et al., 2020b)
based datasets Pull-requests 2 (Sghaier and Sahraoui, 2023) (Zhang et al., 2020b)
Combined datasets Source code and comments 8 18 (Gu et al., 2022) (Lajkó et al., 2022) (Niu et al., 2022) (Prenner and Robbes, 2021) (Saberi et al., 2023b) (Tufano et al., 2022) (Zhang et al., 2022c) (Zhang et al., 2023e)
Programming tasks and test suites/cases 4 (Liu et al., 2023c) (Olausson et al., 2023) (Ouyang et al., 2023) (Shirafuji et al., 2023)
Binary code and related annotations 1 (Al-Kaswan et al., 2023)
Failing test code and error messages 1 (Xia and Zhang, 2023b)
Source code and Q&A pairs 1 (Salza et al., 2022)
@@
Source code and test suites/cases 1 (Tufano et al., 2023)
Source code, methods, and 1 (Li et al., 2023a)
logging statements
Source code, description, and 1 (Liu et al., 2023a)
code environment
@@
Data types play a pivotal role in shaping the architecture and selection of LLMs, as they directly influence the extraction of implicit features and subsequent model decisions(Chan et al., 2023; Ghadhab et al., 2021; Yang et al., 2023c; Shi et al., 2022). The choice of data types can significantly impact the overall performance and generalization ability of the LLMs. We examine and classify the types of SE datasets employed in LLM4SE studies. By investigating the relationship between data types, model architectures, and performance, we seek to shed light on the critical role of data types in the success of LLM4SE applications.
@@
Data type categorization. We classified the data types of all datasets into five categories: code-based, text-based, graph-based, software repository-based, and combined data types. Table 6 describes the specific data included in the data types corresponding to the datasets we summarized from the 229 studies. We can find that most of the studies used text-based datasets, accounting for a total of 104. The dominance of text-based datasets in training LLMs for SE tasks highlights the models’ exceptional natural language processing capabilities. These LLMs excel in understanding and processing textual data, making them an ideal choice for tasks that involve code comprehension, bug fixing, code generation, and other text-oriented SE challenges. Their ability to process and learn from vast amounts of text data enables them to provide powerful insights and solutions for various SE applications. Text-based datasets with a large number of prompts (28) are commonly used in training LLMs for SE tasks to guide their behavior effectively. While understanding the training data may not be essential for closed-source LLMs like ChatGPT, insights into the data handling techniques of other models remain valuable. This is particularly true as black-box models can be fine-tuned with small-sized data inputs during usage. Among the 229 surveyed papers, this understanding is reinforced by the fact that text-based datasets with a large number of prompts are the most frequently used data types for training LLMs in SE tasks. programming problems (14) are also essential as they provide diverse and challenging tasks, allowing models to generalize knowledge and skills for various SE challenges. This combination helps the models develop a robust understanding of software concepts and perform well in a wide range of tasks. There are also SO (i.e., Stack Overflow) posts (9), bug reports (9), programming tasks (and solutions) (7), etc., which are among the more numerous data types in text-based datasets.
@@
The predominance of source code (44) as the most abundant data type in code-based datasets can be attributed to its fundamental role in SE. Source code serves as the foundation of any software project, containing the logic and instructions that define the program’s behavior. Therefore, having a large volume of source code data is crucial for training LLMs to understand the intricacies of software development, enabling them to effectively generate, analyze, and comprehend code in various SE tasks. There are also common data types such as bugs (4) and patches (4) for program repair tasks. Graph-based datasets are used in some research studies for SE tasks, e.g., Kolthoff et al. (Kolthoff et al., 2023) used a dataset composed of screenshots from Google Play Android applications to construct a graphical user interface (GUI) repository in their study on LLM for the rapid prototyping task. These datasets represent code using graph structures, capturing relationships and dependencies between code components.
@@
Software repository-based datasets usually mean data collected from software version control systems (e.g., Git) and issue tracking systems (e.g., GitHub, Jira, etc.). This data includes issues and commits (3), pull requests (2), and so on. The data in software repositories can provide a wealth of information covering all aspects of the software development process, including code evolution history, records of issue fixes and feature improvements, code quality assessments, and so on. These data are valuable for studying behaviors and trends in the software development process, improving software quality and development efficiency, and evaluating the performance of software engineering techniques. Therefore, many studies have used software repository-based datasets for empirical analysis and model training.
@@
Some studies employed combined datasets containing multiple datatypes. Among them, the most common type is “source code and comments”. For instance, Tufano et al. (Tufano et al., 2022) used a dataset comprising source code and comments to train a model and showed that a pre-trained text-to-text converter (T5) model outperforms a previous deep learning model in automating the code review task. Other combinations of data types include “binary code and related annotations”, “failing test code and error messages”, “source code and Q&A pairs”, “source code, description, and code environment”, etc.

### v6

@@ 4. RQ2: How are SE-related datasets collected, preprocessed, and used in LLMs?
Data plays a crucial role in the model training phase (Sun et al., 2022). First, data is collected to obtain diversity and richness to ensure that the model can cope with different scenarios and situations. Second, data is classified to clarify the training objectives of the model and avoid confusion and misinformation. The preprocessing of data is indispensable to clean and transform the data to improve its quality. Finally, data is formatted into a structure suitable for model processing, allowing the LLM to learn the data’s features and patterns effectively. We analyze the reported processes of data collection, data classification, data preprocessing, and data representation in our selected primary studies on LLM4SE.
@@ 4.1. How are the datasets for training LLMs sourced?
Data is an indispensable and critical factor in training LLMs, which determines the generalization ability, effectiveness, and performance of the models (Sun et al., 2022). Adequate, high-quality, and diverse data is critical to allow models to fully learn features and patterns, optimize parameters, and ensure reliability in validation and testing. We first investigate the methods used to obtain the dataset. By analyzing the methods of data collection, we divided the data sources into four categories: open-source datasets, collected datasets, constructed datasets, and industrial datasets. Open-source datasets (Chen et al., 2023c; Khakhar et al., 2023; Wang et al., 2023h; Zeng et al., 2022) refer to publicly accessible collections of data that are often disseminated through open-source platforms or repositories. For example, datasets like HumanEval (Chen et al., 2021b), which consists of 164 manually crafted Python problems, each accompanied by its respective unit tests. The open-source nature of these datasets ensures their credibility and allows for community-driven updates, making them a reliable resource for academic research. Collected datasets (Huang et al., 2018; Tian et al., 2023b; Sghaier and Sahraoui, 2023; Mastropaolo et al., 2022b) are those that researchers compile directly from a multitude of sources, including but not limited to, major websites, forums, blogs, and social media platforms. For instance, researchers (Chan et al., 2023; Salza et al., 2022; Weyssow et al., 2023b; Yang et al., 2022c) often scrape data from Stack Overflow (Overflow, 2023) threads or GitHub (Github, 2023) issues comments to create a dataset tailored to their specific research questions. Constructed datasets (Ezzini et al., 2022; Koide et al., 2023; Kang et al., 2022; Zhang et al., 2022a) are specialized datasets that researchers create by modifying or augmenting collected datasets to better align with their specific research objectives. These modifications can be carried out through manual or semi-automatic methods and may include the generation of domain-specific test sets, annotated datasets, or synthetic data. For example, researchers often take a collected dataset of code snippets and manually annotate them with bug types to create a constructed dataset for studying automated program repair techniques (Fan et al., 2023a; Jin et al., 2023b; Wu et al., 2023a). Industrial datasets (Alhamed and Storer, 2022; Moharil and Sharma, 2022; Wang et al., 2020c) are those obtained from commercial or industrial entities and often contain proprietary business data, user behavior logs, and other sensitive information. These datasets are particularly valuable for research that aims to address real-world business scenarios. However, the acquisition of such datasets is often complicated by issues related to business confidentiality and data privacy. For example, in a collaborative effort with China Merchants Bank (CMB), Wang et al. (Wang et al., 2020c) were able to access 21 projects from CMB’s repositories. Access to such data would likely require non-disclosure agreements and other legal safeguards to protect business interests. Each of these dataset types offers unique advantages and challenges, and the choice between them should be guided by the specific requirements and constraints of the research project at hand.
Figure 6. The collection strategies of LLM-related datasets.
@@
Fig. 6 shows the collection strategies of LLM-related datasets. As can be seen from the data in the figure, 235 studies used open-source datasets for training LLMs. One of the main reasons for using open-source datasets in LLM training is their authenticity and credibility. Open-source datasets usually contain real-world data collected from various sources (such as relevant studies that have been conducted), which makes them highly reliable and representative of real-world scenarios. This helps LLMs learn from real examples to better understand real-world applications and improve their performance. Second, since LLMs are a topic that has just recently emerged, a lack of suitable training sets does exist. Therefore, researchers often collect data from sites such as Stack Overflow and GitHub and build datasets to make the data more composite for SE tasks. Among the 395 papers we studied, we discovered that merely six studies utilized industrial datasets. This suggests a potential misalignment between the properties of datasets used in academic research and those encountered in real-world industrial contexts. This divergence underscores the need for future research to investigate industrial datasets, thereby ensuring that LLMs are applicable and robust across both academic and industrial scenarios.
@@ 4.2. What types of SE datasets have been used in existing LLM4SE studies?
Category Data type Total
Text-based Programming tasks/problems (42) Prompts (33) 151
datasets SO (i.e. Stack Overflow) posts (12) Bug reports (11)
Requirements documentation (9) APIs/API documentation (8)
Q&A pairs (6) Vulnerability descriptions (4)
Reviews (4) Logs (3)
Methods (3) Project issues (3)
Code comments (2) Theorems (2)
Buggy text (1) Dockerfiles (1)
Outage descriptions (1) Semantic merge conflicts (1)
Site text (1) Software development tasks (1)
User intents (1) Software specifications (1)
User reviews (1)
Code-based Source code (60) Bugs/Buggy code (16) 103
datasets Vulnerable source code (8) Patches (4)
Code changes (3) Test suites/cases (3)
Bug-fix pairs (2) Error code (2)
Error-fix pairs (1) Flaky test cases (1)
Identifiers (1) Labeled clone pairs (1)
Packages (1)
Graph-based GUI Images (1) 1
datasets
Software Code repository (9) Android apps (3) 20
repository Issues and commits (3) Pull-requests (2)
-based datasets Industrial projects (1) Open-source projects (1)
Web applications (1)
@@
Combined Programming tasks and test suites/cases (17) Source code and comments (12) 55
datasets Programming tasks and solutions (8) Source code and description (3)
Code-text pairs (2) Souce code and API usage sequences (2)
Source code and test suites/cases (2) Bug report and test suites/cases (1)
Buggy code and comments (1) Buggy code and solutions (1)
Code files and summaries (1) Binary code and related annotations (1)
Failing test code and error messages (1) Source code and Q&A pairs (1)
Source code, methods, and logs (1) Vulnerable code and description (1)
  • *See Appendix A for the full table including references.
@@
Data types play a pivotal role in shaping the architecture and selection of LLMs, as they directly influence the extraction of implicit features and subsequent model decisions(Chan et al., 2023; Ghadhab et al., 2021; Yang et al., 2023f; Shi et al., 2022). The choice of data types can significantly impact the overall performance and generalization ability of the LLMs. We examine and classify the types of SE datasets employed in LLM4SE studies. By investigating the relationship between data types, model architectures, and performance, we seek to shed light on the critical role of data types in the success of LLM4SE applications.
@@
Data type categorization. We classified the data types of all datasets into five categories: code-based, text-based, graph-based, software repository-based, and combined data types. Table 7 describes the specific data included in the data types corresponding to the datasets we summarized from the 395 studies. We can find that most of the studies used text-based datasets, accounting for a total of 151. The dominance of text-based datasets in training LLMs for SE tasks highlights the models’ exceptional natural language processing capabilities. These LLMs excel in understanding and processing textual data, making them an ideal choice for tasks that involve code comprehension, bug fixing, code generation, and other text-oriented SE challenges. Their ability to process and learn from vast amounts of text data enables them to provide powerful insights and solutions for various SE applications.
The most prevalent type of data utilized in training LLMs for SE tasks is programming tasks/problems with 42 instances observed among the surveyed papers. This dominance can be attributed to the diverse and challenging nature of programming problems, which provide LLMs with opportunities to generalize knowledge and skills across various SE challenges, fostering a robust understanding of software concepts and enhancing performance across a wide range of tasks, including code generation, code completion, and code summarization, etc. Prompts follow closely behind programming tasks, with 33 instances observed in the surveyed papers, providing task-specific guidance to LLMs, serving as cues or instructions for the models, and helping them understand the context and requirements of SE tasks. This combination helps the models develop a robust understanding of software concepts and perform well in a wide range of tasks. There are also SO (i.e., Stack Overflow) posts (12), bug reports (11), etc., which are among the more numerous data types in text-based datasets.
@@
The predominance of source code (60) as the most abundant data type in code-based datasets can be attributed to its fundamental role in SE. Source code serves as the foundation of any software project, containing the logic and instructions that define the program’s behavior. Therefore, having a large volume of source code data is crucial for training LLMs to understand the intricacies of software development, enabling them to effectively generate, analyze, and comprehend code in various SE tasks. There are also common data types, such as bugs/buggy code (16) and patches (4), for program repair tasks. Additionally, vulnerable source code (8) is used for vulnerability detection tasks. Graph-based datasets are used in some research studies for SE tasks, e.g., Kolthoff et al. (Kolthoff et al., 2023) used a dataset composed of screenshots from Google Play Android applications to construct a graphical user interface (GUI) repository in their study on LLM for the rapid prototyping task. These datasets represent code using graph structures, capturing relationships and dependencies between code components.
@@
Software repository-based datasets are compilations of data extracted from version control systems, such as Git repositories, containing code, documentation, and related artifacts. This data includes Code repository (3), issues and commits (3), and so on. The data in software repositories can provide a wealth of information covering all aspects of the software development process, including code evolution history, records of issue fixes and feature improvements, code quality assessments, and so on. These data are valuable for studying behaviors and trends in the software development process, improving software quality and development efficiency, and evaluating the performance of software engineering techniques. Therefore, many studies have used software repository-based datasets for empirical analysis and model training.
@@
Some studies employed combined datasets containing multiple datatypes. Among them, the most common type is “programming tasks and test suites/cases”. Other combinations of data types include “source code and comments”, “programming tasks and solutions”, “source code and description ”, “code-text pairs”, etc.